OpenStudy (anonymous):

how to simplify the square root of 18 to it's radical form?

OpenStudy (ybarrap):

Factor out the 18 under the radical and simplify: $$ \sqrt{18}=\sqrt{3\times2\times3}=\sqrt{3^2\times2}=\sqrt{3^2}\times\sqrt{2} $$ What's next?

OpenStudy (anonymous):

is it \[3 \sqrt{2}?\]

OpenStudy (sweetburger):

yes

OpenStudy (anonymous):

ok cool, thanks guys